Question 37
ProbabilityThe probabilities that John and Jane will pass an examination are 0.9 and 0.7 respectively. Find the probability that at least one of them will pass the examination.
- A.0.28
- B.0.67
- C.0.72
- D.0.97
Show the answer and explanation
Answer: D (0.97)
Explanation
P(neither) = 0.1 × 0.3 = 0.03, so P(at least one) = 0.97.